Operation Summary for May 19, 2010 - May 21, 2010 |
Wednesday May 19 The day shift began with Operations monitoring the Tevatron (TeV) store 7820 (89.96E30) and the Recycler stash (114.0E10), with the Antiproton Source (Pbar) stacking (11.0E10), with the Linac, Booster, and the Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, with beam to the Fermilab Test Beam Facility (FTBF) T-1005 experiment, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, Operations terminated store 7820 at 10 AM. Operators prepared CDF, DZero, and the TeV for a short round of accesses, which lasted until 3:30 PM. The CDF collision hall interlocks dropped and interlock technicians investigated. They found no obvious hardware failures. Operators resecured the area at 1:45 PM. During the access, cryo system personnel replaced the flywheel for the TeV sector A4 wet engine. An AP1 line quadrupole power supply (Q107) failed to ramp. Pbar personnel configured the supply for DC operation and continued to run until they had a chance to investigate its regulation problem. Operations halted stacking at 1 PM for 30 minutes so Pbar personnel could repair the power supply. Booster and Linac personnel took the opportunity to look at a Booster RF station (BRF18) and to make an I- Source dome access. During the I- dome access, communication with the controls system was disrupted. Linac personnel switched to the H- Source shortly thereafter. The Pelletron’s blower motor failed; experts accessed the MI-31 enclosure to investigate. Operations put the TeV through a dry squeeze. During the evening shift, operators discovered a TeV dipole power supply (HE48) that had tripped off. TeV personnel reseated the load connections and the supply came on. Another dry squeeze and TeV orbit check showed no further problems. Operations put the TeV into HEP shot setup at 8 PM, but that ended abruptly with magnet quenches at C4, F4, and D0 during the low beta squeeze. Operators discovered that HE48 had tripped off during acceleration. Operations disabled the T-1005 beam at 6:18 PM. Mechanical Support personnel had found and replaced the blower motor controller for the Pelletron. The C48 Kautzky valve showed a leak indication following the quench. Operators prepared the TeV A-E enclosures for access, but the leak indication cleared before any accesses were made. CHL suggested Operations continue with the quench recovery and TeV turn on. Thursday May 20 The midnight shift began with the TeV turn on just starting, with Operations monitoring the stash (106.8E10), with Pbar stacking (15.0E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations put the TeV through a dry squeeze and orbit check. Operations put the TeV into shot setup at 3:30 AM and established store 7826 at 4:20 AM with an average initial luminosity (AIL) of 207E30. An AP1 line quadrupole power supply (Q105) tripped off on phase imbalance at 4 AM. Operators called in Pbar personnel to make repairs; they replaced a SCR gate firing board. Operations established beam to T-1005 at 5:10 AM. Operations halted the NuMI beam at 5:30 AM and prepared MI-65 for a scheduled access. Pbar resumed stacking at 5:45 AM. During the day shift, Linac personnel adjusted the beam pulse width. Pelletron personnel accessed MI-31 to reverse the polarity of the blower motor. An industrial water leak was found near the MI-8 service building. NuMI personnel accessed MI-65 to check their humidifiers; they were out of access and had resumed taking beam by 12:15 PM. During the evening shift, Operations disabled beam to T-1005 at 6:04 PM. Operations terminated store 7826 at 11 PM and put the TeV into shot setup. Operators checked the Switchyard enclosures C, D, and E, and the F1 manhole for LCW leaks - none were found. Friday May 21 The midnight shift began with the TeV in shot setup, with Pbar stacking (41.0E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established store 7827 at 12:20 AM with an AIL of 374E30. Operations established beam to t-1005 at 4:57 AM. Operations reported a minimal amount of downtime due to problems with the Booster and Linac RF. The Plans for Friday and the Weekend The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow Booster personnel to bring BRF18 back online and for Controls personnel to replace a MI-8 crate power supply. |
|
More Information Comments and Suggestions |
| last modified 05/21/2010 email Fermilab |
| Security, Privacy, Legal | |